OpenStudy (anonymous):

use the first derivative test to find any relative extrama of the function. g(x)=(3/2)sin((pi*x)/2-1), [0,4]

OpenStudy (anonymous):

\[g(x) =\frac{ 3 }{ 2 }\sin (\frac{ pix }{ 2 }-1)\]